

********************************************************************
UARS WEEKLY STATUS REPORT - 06 February 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occultation
scan data during this past week without apparent problems.
***********
PEM
***********
PEM is currently off.
PEM is awaiting evaluation by GSFC of its proposed activation procedure.
The minimal configuration for PEM AUX activation is the zenith VMAG, MEPS1,
MEPS2, and HEPS1 units, only. This configuration is driven by the capabilities
of the spacecraft OBC and may increase if OBC program area increases. PEM
activation is proposed to be OBC driven when there exists plenty of power to
operate PEM. It is anticipated that the OBC will activate PEM on the daylight
portions of every orbit.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
MLS was reactivated successfully on Jan 30, in the reduced 85 watt power mode.
MLS instrument operations have continued flawlessly. Full vertical scans have
been performed on 6 of the last 7 days. MLS is scheduled to operate through
Feb 12.
***********
HRDI
***********
HRDI is currently in low-power survival mode. Our next scheduled turnon is
Feb. 10.
***********
WINDII
***********
WINDII was activated on February 3 and is performing normally.
Observations are planned to continue until February 12.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF has received a total of 367 CD-ROM requests as of 06-February-1998.
There were two orders received and two orders shipped.
The CDHF coordinated the ingest of the new SSPP Calibration file with the FOT,
FDF and Solstice. PI representatives were notified when the file
(CAL_SSSPPCALBR_D0004.XABMATRICES_V0005_C01_PROD) was available.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
25 days PEM (3.32) PROD (1379R-879C=500L)
125 days MLS (4.90) PROD (1344R-323C-3F=1021L)
0 day CLAES (7.0) ENG (17R-0C=17L)
1 day SOL (4.52) PROD (1R-1C=0L)
1 DAY OBC (3.4) PROD (1R-1C=0L)
The CDHF continues to work with NSI to develop a support plan for the
entire UARSNET network, focusing on Instrument RACs first, redesign issues,
and then the global user community.
Pending software upgrades: IDL 5.0, OpenIngres.
***********
FOT
***********
REPORTING PERIOD
This report covers 24 January 1998 (Orbit 34807, GMT Day 024) through 30
January 1998 (Orbit 34910, GMT Day 030).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HRDI 028/2312 34881 Off
PEM 029/0055 34882 ZEPS on
PEM 029/0656 34885 AXIS on
HALOE 029/1026 34888 Off
MLS 030/0434 34899 On
A summary of Instrument operations supplied by the Mission Planning Group
for the month of January 1998 is contained in Appendix F, UARS Instrument
Operations Summary.
The beta angle is now increasing toward a maximum angle of 72.6 degrees on
05 February 1998 (DOY 036).
A series of SSPP gridded offset maneuvers requested by SUSIM was performed
on GMT days 024 through Day 029. The offsets utilized the 5 and 10
arcminute intervals, as opposed to the 15 and 30 arcminute intervals.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 24.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.3 degrees C. Additional battery performance data for
each battery on 24 January 1998 and 30 January 1998 are provided in
Appendix D.
The clock error ranged from -10.3 to +10.8 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = none
% OUT-OF-LIMITS = none
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 03
% HALOE 01
% HRDI 01
% MLS 01
% PEM 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 02
TRANSPONDER FREQUENCIES:
Transponder A: 98/020 - 2287.497001
Transponder B: 98/021 - 2287.498836
GROUND SYSTEM OPERATION
One (1) operational support problem(s) occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/029 show the observatory data loss to be
20 hours, 41 minutes, 48.244 seconds (an increase of 12.928 seconds since
last reported). This is a 0.0371 percent data loss which equals a 99.9629
percent data capture for the mission.
The increase of 12.928 seconds occurred over a five day period as follows:
7 01.568 seconds on Day 023 due to insufficient overlap,
7 00.448 seconds on Day 024 due to insufficient overlap,
7 07.488 seconds on Day 026 due to insufficient overlap,
7 02.432 seconds on Day 027 due to insufficient overlap, and
7 00.992 seconds on Day 029 due to insufficient overlap.
No testing of the new CMS software release (R7B1V2) occurred this week.
UARS began testing IP return-link telemetry this week. The FOT has been
utilizing a RparallelS AP configuration to monitor the IP data, which comes
into MSOCC on the MS440 (I-channel) and MS442 (Q-channel) lines. Thus far,
the FOT has verified successful I-channel telemetry, but has been unable to
obtain Q-channel OBC dump data. The MSOCC NASA System Engineer is aware of
the Q-channel problem and is working with NASCOM to rectify this situation.
Operational processing of level-0 telemetry analysis files continues in
parallel on both the UMPGAL (new MPG Alpha machine) and CDHF computer
systems. No level-0 file transfer failures (between CDHF and UMPGAL)
occurred this week.
SUSIM operations submitted a request for the FOT to schedule six additional
SOLSTICE offsets for the week of 01 February 1998. One offset will take
place each day, beginning with GMT Day 031 and ending with GMT Day 036. In
addition, SUSIM has also submitted a request for similar offsets to occur
on GMT Days 038-043.
OTHER SIGNIFICANT EVENTS
Work on the upgrade of the OBC Power Monitor (PMON) software continued. A
second iteration of redesign and coding is nearly completed. A test image
should be available by late this week or early next week. The test version
of the flight software will be designated version 62, revision D (V62D).
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current
status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
Dimitrios Mantziaras passed the Operations Support Engineer (OSE)
certification test on 27 January 1998. Dimitrios began his regular shift
rotation (Deryck ChetramUs crew) on 26 January 1998.
MEETINGS
UARS attended a meeting on 28 January 1998 devoted to the Year 2000 (Y2k)
impact on operational elements specifically the UARS Test and Training
Simulator (UTTS). The UTTS software personnel feel that the simulator will
not be a problem. Due to the "age" of the operating system and layered
products, the plan is that the system time on the VAX will be set to a
"wrong" date. There's no problem with this because the simulator does not
use the actual time for anything. The simulator does, however, expect the
UARS TAC to have the correct time. The UARS POCC, including the TAC, RUPS,
and APs, is a much bigger issue.
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 25 February
1998.
UARS Flight Operations plans to attend the Program Experience portion of
the Space Power Workshop at JPL in April 1998. UARS FOT plans on
presenting an overview of the UARS batteries and their performance.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust
will probably be required in 1999.
FACILITIES
no change
PERSONNEL CHANGES
no change
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 30 January 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
34818 024/1939 98-030 OBC Clk Adj fr 0.00 to 20.6 msec/day
34854 027/0450 98-031 OBC Clk Adj fr 20.6 to 0.00 msec/day
34875 028/1452 98-032 OBC Clk Adj fr 0.00 to 20.6 msec/day
34881 028/2312 98-033 HRDI Power Off
34882 029/0055 98-034 PEM ZEPS On
34885 029/0656 98-035 PEM AXIS On
34888 029/1026 98-036 HALOE Power Off
34899 030/0434 98-037 MLS Power On
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 30 January 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
none this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
none this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 30 January 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
34818 024/19:48 (24 Jan 98) from 00.0 to 20.6 msec/day
34854 027/04:56 (27 Jan 98) from 20.6 to 00.0 msec/day
34875 028/14:55 (28 Jan 98) from 00.0 to 20.6 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 30 January 1998
24 Jan 1998 - GMT Day 024 Beta = 32.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.6 0.000 0.0
2 +89.6/ -5.6 +6.44/+4.83 +31.2/-16.0 25.1 1.025 22.5
3 +61.6/ -11.2 +5.15/+3.87 +30.4/-16.0 25.1 1.026 22.0
30 Jan 1998 - GMT Day 030 Beta = 56.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.68/-4.94 +0.0/ +0.0 18.6 0.000 0.0
2 +11.2/ -67.2 +5.15/+3.23 +26.8/-16.0 24.8 1.021 21.2
3 +56.0/ -11.2 +3.87/+2.29 +25.6/-16.4 24.8 1.024 20.4
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
(no change)
_______________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 30 January 1998
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-004 FOT 34874 028/1232 N/A TAC given to MNX during pass
98-001 FOT 34910 030/2228 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
_______________________________________________________________
APPENDIX F
UARS INSTRUMENT OPERATIONS SUMMARY
Week ending 30 January 1998
INSTRUMENT OPERATIONS SUMMARY INFORMATION FOR 1997
FOURTH QUARTER 1997 SUMMARY (hours)
UNPLANNED UNSCHEDULED
ON OFF OFF OFF
(Insufficient (Other than
Instrument (Actual) (Scheduled) Power) Power Problem) COMMENTS
---------- ------ ----------- ----------- ------------- -----------
-
ACRIM 2038 170 0 0
HALOE 1040 1168 0 0
HRDI 610 1566 32 0
MLS 772 1396 40 0
PEM (axis) 566 1636 6 0
(zeps) 199 1994 15 0
(neps) 0 2208 0 0
SOLSTICE 2208 0 0 0
SUSIM 2038 170 0 0
WINDII 575 1617 16 0
Cold side of spacecraft was facing northward: 44.5 days.
Cold side of spacecraft was facing southward: 47.5 days.
INSTRUMENT OPERATIONS MONTHLY SUMMARY FOR OCTOBER 1997 (hours)
Instrument On Scheduled Off Unscheduled Off
---------- ------ ------------- ---------------
ACRIM 574 170 0
HALOE 475 269 0
HRDI 262 482 0
MLS 257 487 0
PEM (axis) 80 664 0
(zeps) 0 0 0
(neps) 0 0 0
SOLSTICE 744 0 0
SUSIM 574 170 0
WINDII 127 617 0
Cold side of spacecraft was facing northward: 4.5 days.
Cold side of spacecraft was facing southward: 26.5 days.
INSTRUMENT OPERATIONS MONTHLY SUMMARY FOR NOVEMBER 1997 (hours)
UNPLANNED UNSCHEDULED
ON OFF OFF OFF
(Insufficient (Other than
Instrument (Actual) (Scheduled) Power) Power Problem) COMMENTS
---------- ------ ----------- ----------- ------------- -----------
-
ACRIM 720 0 0 0
HALOE 216 504 0 0
HRDI 158 530 32 0
MLS 302 378 40 0
PEM (axis) 186 528 6 0
(zeps) 118 587 15 0
(neps) 0 720 0 0
SOLSTICE 720 0 0 0
SUSIM 720 0 0 0
WINDII 167 553 0 0
Cold side of spacecraft was facing northward: 23.5 days.
Cold side of spacecraft was facing southward: 6.5 days.
INSTRUMENT OPERATIONS MONTHLY SUMMARY FOR DECEMBER 1997 (hours)
UNPLANNED UNSCHEDULED
ON OFF OFF OFF
(Insufficient (Other than
Instrument (Actual) (Scheduled) Power) Power Problem) COMMENTS
---------- ------ ----------- ----------- ------------- -----------
-
ACRIM 744 0 0 0
HALOE 349 395 0 0
HRDI 190 554 0 0
MLS 213 531 0 0
PEM (axis) 300 444 0 0
(zeps) 81 663 0 0
(neps) 0 744 0 0
SOLSTICE 744 0 0 0
SUSIM 744 0 0 0
WINDII 281 447 16 0
Cold side of spacecraft was facing northward: 16.5 days.
Cold side of spacecraft was facing southward: 14.5 days.
INSTRUMENT OPERATIONS SUMMARY INFORMATION FOR 1998)
INSTRUMENT SUMMARY FOR MONTH OF JANUARY 1998 (hours)
UNPLANNED UNSCHEDULED
ON OFF OFF OFF
(Insufficient (Other than
Instrument (Actual) (Scheduled) Power) Power Problem) COMMENTS
---------- ------ ----------- ----------- ------------- -----------
-
ACRIM 744 0 0 0
HALOE 442 302 0 0 (note 1.)
HRDI 213 531 0 0
MLS 120 624 0 0
PEM (axis) 191 553 0 0
(zeps) 213 531 0 0
(neps) 0 744 0 0
SOLSTICE 744 0 0 0
SUSIM 744 0 0 0
WINDII 0 744 0 0
Notes:
1. HALOE instrument team requested turn off of HALOE
14 hours earlier than originally scheduled due to
high TBAT temperature being encountered at the high
solar beta angle.
Cold side of spacecraft was facing northward: 15.5 days.
Cold side of spacecraft was facing southward: 15.5 days.